I need help to solve this problem: Assume that you plan to use a significance level of alpha=0.05 to test the claim that p 1= 2. Use the given sample sizes and numbers of successes to find the ​p-value for the hypothesis test. n 1=​50, n 2=75 x 1=​20, x 2=15 A. 0.0146 B. 0.0001 C. 0.0032 D. 0.1201
Answer #1

Solution :

Given that,

This is the two tailed test .

The null and alternative hypothesis is ,

H0 : P1 = P2

Ha : P1    P2

1 = x1 / n1 = 20 / 50 = 0.4

2 = x2 / n2 = 15 / 75 = 0.2

= (x1 + x2) / (n1 + n2) = (20 + 15) / (50 + 75) = 0.28

1 - = 0.72

Z = (1 - 1) / * (1 - ) (1 / n1 + 1 / n2)

Z = (0.4 - 0.2) / 0.28 * 0.72 (1 / 50 + 1 / 75)

Z = 2.44

P(z > 2.44) = 1 - P(z < 2.44) = 0.0073

P-value = 2 * 0.0073 = 0.0146

A)